OpenStudy (anonymous):

will give medals if somebody can check and see if i chose the right answer Find the non-extraneous solutions of the square root of the quantity of x plus 3 minus 4 equals x minus 1 x = −3 and x = −2 x = −2 x = −3 x = 5<--- i chose this one

OpenStudy (anonymous):

@djcool31

OpenStudy (anonymous):

@DullJackel09 @divu.mkr @Disco619 @angel12310 @androidonyourface

OpenStudy (anonymous):

somebody plz help me

OpenStudy (anonymous):

@clara.university @FlvsGirl @Yuba @Vallarylynn16 @jojo4eva

OpenStudy (anonymous):

@BeginnerCountrySinger @binks @bluebrandon @him1618 @horsecrazy366

OpenStudy (anonymous):

@BeginnerCountrySinger do you think im right?

OpenStudy (anonymous):

@zimmah do you think you can help me

OpenStudy (anonymous):

i'll try

OpenStudy (anonymous):

i just want to double check and see if im right

OpenStudy (anonymous):

so, \[\sqrt{x+3}-4=x-1\] did i read that right?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

ok, it's not 5

OpenStudy (anonymous):

so then it must be -2 and -3 right

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

thanks and do you think you can help me with one more question

OpenStudy (anonymous):

most likely i can, i'll try at least

OpenStudy (anonymous):

Simplify 12 over the quantity of negative 3 plus 3i

OpenStudy (anonymous):

i'm not sure how to interpret that, can you draw it?

OpenStudy (anonymous):

do you mean \[\frac{ 12 }{ -3+3i }\]

OpenStudy (anonymous):

or \[\frac{ 12 }{ -3 }+3i\]

OpenStudy (anonymous):

or something else?

OpenStudy (anonymous):

the problem looks like this 8 over 2+2i

OpenStudy (anonymous):

i'm not sure what you mean by over

OpenStudy (beginnercountrysinger):

i'm only in geometry hun .-.

OpenStudy (anonymous):